OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? g(x) − f(x) = 3^x − 8 − 5x g(x) − f(x) = 3^x − 10 − 5x g(x) − f(x) = −2x − 8 − 3^x g(x) − f(x) = −2x − 10 − 3^x

OpenStudy (anonymous):

Please help :(

OpenStudy (bradely):

g(x) − f(x) =(3^x − 9)-(5x − 1 ) =3^x-9-5x+1 =3^x-5x-9+1 =3^x-5x-8 Option 1 is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

\[g(x) -f(x) = 3x -9 -(5x -1) => 3x -9 -5x + 1 = -2x -8\] 3rd option I guess